OpenStudy (madeleine2):

Create an algebraic expression for each of the following special products patterns. Use the special products formulas to complete the multiplication. Show all work for full credit. A. perfect square trinomial formula B. difference of perfect squares formula

OpenStudy (will.h):

A: Any trinomial of the form ax^2 + bx + c is said to a perfect square if it satisfies the condition b^2 = 4ac Example: \[(ax)^2 + 2abx+ b^2 = (ax + b)^2\]

OpenStudy (will.h):

B: \[x^2 - 2xy + y^2 = (x - y)^2\]